An architect designs an rectangular flower garden such that width is exactly two-thirds of the lenght. If 290 feet of antique picket fencing are to be used to enclose thw garden, find the dimension of the garden.
Answer by jorel1380(3719)

2l+2(2/3)l=290
2l+4/3l=290
6l+4l=870
10l=870
l=87
w=58
The garden was 87 x 58 feet..